### v1

# Analyzing Message-Code Inconsistency in AI Coding Agent-Authored Pull Requests

Abstract.
[…]
LLMs, Code Review, Human-AI Collaboration, AI Trustworthiness, Empirical SE, Mining Software Repositories, LLM4Code, AI4SE
$c$s: Software and its engineering Collaboration in software development
1. Introduction
AI coding agents are increasingly acting as autonomous teammates in software development (Watanabe et al., 2025; Wang et al., 2025b; Gong et al., 2025a), generating code and opening pull requests (PRs) with natural-language descriptions that communicate intent (Ogenrwot and Businge, 2025). Unlike human-written PRs, AI agent-authored PRs (Agentic-PRs) are largely black boxes (Watanabe et al., 2025), forcing reviewers to rely on PR descriptions to understand changes and assess correctness (Zhang et al., 2022; Ford et al., 2019). As a result, the reliability of PR descriptions is critical for enabling effective human-AI collaboration.
However, AI-generated descriptions (title + body) are not always faithful to the underlying code, as generative models can produce hallucinated or incorrect statements (Huang et al., 2025b; Twist et al., 2025), undermining trust in AI agents. While message-code inconsistency (MCI) has been studied for human-written commit messages (Dong et al., 2023; Zhang et al., 2025; Wen et al., 2019), little is known about how often Agentic-PR descriptions misalign with code changes, what inconsistency types occur, or whether such misalignment affects reviewer trust and PR outcomes.
To address this gap, we conducted an empirical study analyzing 23,247 Agentic-PRs from the AIDev dataset (Li et al., 2025) using PR message-code inconsistency (PR-MCI) and made the following contributions:
  1. (1)
    Annotated dataset: We release 974 manually annotated PRs (432 partial/misaligned) drawn from AIDev to support the development of more reliable AI coding agents.
  2. (2)
    Prevalence (RQ1): We found 406 high-MCI PRs (1.7%) with a 20-fold variation across agents, suggesting that a non-trivial fraction of PRs could mislead reviewers.
  3. (3)
    Taxonomy (RQ2): We identified eight PR-MCI types, finding that Phantom Changes (descriptions claim unimplemented changes) dominated (45.4%).
  4. (4)
[…]
Overall, our findings indicate that while AI agents can generate PRs at scale, the reliability of their descriptions varies substantially and is associated with reviewer trust and PR outcomes, offering actionable insights for developers, tool builders, and researchers.
2. Background and Related Work
Agentic-PRs. AI coding agents are increasingly capable of autonomously authoring and submitting PRs in real-world software projects (Ogenrwot and Businge, 2025; Brookes et al., 2025; Watanabe et al., 2025). In these Agentic-PRs, natural-language descriptions play a critical role in communicating intent and scope to human reviewers. However, generative AI systems are known to produce hallucinated or incorrect statements (Huang et al., 2025b; Twist et al., 2025), raising concerns about the reliability and trustworthiness of Agentic-PRs.
Message-Code Inconsistency. Commit messages and PR descriptions serve as documentation of developer intent. When these descriptions misalign with actual code changes, they create message-code inconsistency (MCI) (Wang et al., 2025a; Dong et al., 2023). Recent work has studied MCI in human-written commit messages and benchmarked LLMs for detecting text-code inconsistencies (Dong et al., 2023; Zhang et al., 2025), highlighting challenges in generating faithful change descriptions.
Gap in PR-MCI. While MCI has been studied for human-written commit messages (Dong et al., 2023; Zhang et al., 2025), and prior work has analyzed PR text for quality concerns (Karmakar et al., 2022), very few prior work examines inconsistency in Agentic-PR descriptions. This gap is critical because agents generate descriptions automatically without human oversight, potentially producing misleading explanations at scale. PR decision-making research shows that description quality is associated with acceptance rates (Zhang et al., 2022), and automatically created PRs can differ in how maintainers interact with them (Wyrich et al., 2021). Our work extends this research by examining how description-code consistency is associated with outcomes for Agentic-PRs. Note that PR descriptions differ from commit messages: PR descriptions may span multiple commits and serve as the primary communication channel for reviewers.
3. Methodology and Experiment Setup
Figure 1. Methodology workflow for PR-MCI analysis.
We structured the study around three research questions (RQs):
  • RQ1 (Prevalence): How frequently do Agentic-PR descriptions misalign with their code changes?
  • RQ2 (Taxonomy): What types of message-code inconsistencies occur in AI-generated PRs?
  • RQ3 (Impacts): Is message-code inconsistency associated with PR acceptance and review effort?
Figure 1 illustrates our workflow. We used the AIDev dataset (Li et al., 2025) and analyzed the AIDev-pop subset (33,596 PRs from 2,807 repositories with >>100 stars) via the pull_request split. After filtering for closed PRs, permissive licenses (MIT or Apache 2.0), and the six most common task types (95.2%), the final dataset contains 23,247 PRs authored by five AI agents.
3.1. Measuring PR-MCI (RQ1)
Following prior MCI research (Wen et al., 2019; Zhang et al., 2025), we used PR message-code inconsistency (PR-MCI) as the degree to which a PR description (title + body) diverges from the underlying code changes. We measured PR-MCI using a heuristic similarity score that combines three complementary signals: scope adequacy $s_{s}$ (whether description verbosity matches code churn), file-type consistency $s_{f}$ (whether mentioned file types, e.g., tests or documentation, are actually modified), and task-type alignment $s_{t}$ (whether description language matches the labeled task type). The similarity score $s\in[0,1]$ aggregates these signals using a weighted scheme (Kaliszewski and Podkopaev, 2016):
(1) $s=0.3\cdot s_{s}+0.4\cdot s_{f}+0.3\cdot s_{t}$
where $s_{f}$ receives the highest weight (40%) for providing the most concrete and verifiable evidence of misalignment (e.g., claiming test changes when no test files are modified), while $s_{s}$ and $s_{t}$ each receive 30% as complementary but less definitive signals (Barnett et al., 2016).
To validate the PR-MCI score and calibrate the decision threshold, two authors manually annotated 600 PRs, stratified by agent, task type, and score bins, as aligned, partially aligned, or misaligned, achieving strong agreement ($\kappa=0.892$). PRs with similarity below $\theta=0.61$ are labeled high-MCI, where $\theta$ is selected by optimizing F1 on labeled validation data (Fan and Lin, 2007) after conducting sensitivity analysis across threshold ranges (Lipton et al., 2014); five-fold cross-validation shows stable threshold selection (mean=0.606, std=0.008) (Zhang et al., 2025). We also evaluated alternative methods: an embedding model (Qwen3-Embedding-0.6B (Qwen Team, 2024)) performs poorly (F1=0.150), and an agreement method (both methods must agree) achieves lower F1 (0.567 vs. 0.630), indicating both are insufficient (see Table 1).
Consistent with large-scale inconsistency studies (Wen et al., 2019), we computed PR-MCI automatically for all PRs and reported prevalence with 95% Wilson confidence intervals.
3.2. PR-MCI Taxonomy Development (RQ2)
Following mixed-method inconsistency studies that derive taxonomies from manually coded, automatically retrieved candidates (Wen et al., 2019), we developed a PR-MCI taxonomy through an iterative process: GPT-5.2 (OpenAI, 2025) first generates an initial codebook with eight categories, which two annotators refine through discussion before manually classifying 432 partial/misaligned PRs from 974 PRs (600 validation + 374 additional high-MCI PRs from RQ1).
Note that RQ1 and RQ3 use a strict binary definition of high-MCI (treating partially aligned PRs as aligned) to identify only clearly misaligned PRs, whereas RQ2 includes both partial and misaligned PRs to capture a broader range of inconsistency patterns for a more comprehensive taxonomy. As in prior work, the taxonomy reflects the analyzed subset and may not capture all inconsistency types.
3.3. Statistical Tests (RQ3)
To examine associations between PR-MCI and outcomes, we compared high- and low-MCI PRs using standard statistical tests. Following PR decision research (Zhang et al., 2022), we analyzed acceptance rate, review count, comment count, and time to merge. We applied Chi-square tests for binary outcomes and Mann-Whitney U tests for skewed continuous variables, reported Cramér’s V and Cliff’s $\delta$ as effect sizes. Significance is assessed at $p<0.05$.
To control for confounders, we fitted regression models including log-transformed code churn, files changed, task type, and agent, following established PR outcome research (Zhang et al., 2022).
4. Results
4.1. RQ1: Prevalence and Patterns of PR-MCI
Table 1. High-MCI Prevalence and Validation Metrics. [lower, upper] show 95% Wilson confidence intervals. Highlighted values are from the primary Heuristic MCI scoring method.
Category Heuristic (PRIMARY) Embedding Agreement Overall Prevalence 1.7% [1.6, 1.9] 60.0% 1.4% High-MCI Prevalence by Agent (%) GitHub Copilot 8.7% [7.7, 9.9] 83.5% 7.1% Cursor 4.5% [3.2, 6.4] 76.8% 3.4% Claude Code 0.9% [0.2, 3.3] 74.5% 0.9% OpenAI Codex 0.8% [0.6, 0.9] 52.7% 0.6% Devin 0.4% [0.2, 0.7] 75.5% 0.2% High-MCI Prevalence by Task Type (%) Chore 4.0% [2.7, 5.9] 54.5% 3.0% Refactor 3.5% [2.7, 4.5] 47.7% 2.2% Bug Fix 2.1% [1.8, 2.5] 49.5% 1.6% Feature 1.5% [1.3, 1.8] 69.7% 1.3% Documentation 1.0% [0.7, 1.5] 57.4% 0.8% Test 1.0% [0.6, 1.5] 49.3% 0.6% Validation Metrics (n=600) Precision 0.742 0.083 0.760 Recall 0.548 0.786 0.452 F1 0.630 0.150 0.567
[…]
Table 1 reports validation metrics and prevalence by agent and task type. Overall, 406 out of 23,247 Agentic-PRs (1.7%) exhibited high PR-MCI (similarity score below 0.61), where descriptions poorly match code changes, suggesting that a small but non-trivial fraction could mislead reviewers. The prevalence revealed a 20-fold difference across agents: GitHub Copilot had the highest rate (8.7%, 234 high-MCI PRs out of 2,675), followed by Cursor (4.5%, 31 out of 682). Sensitivity analysis (threshold 0.55-0.65) suggested that relative ordering remained consistent across thresholds  [^1] Sensitivity analysis results are available in our replication package..
[…]
RQ1: 406 PRs (1.7%) exhibited high PR-MCI, suggesting that a non-trivial fraction of PRs could mislead reviewers. In particular, GitHub Copilot showed 8.7% (20x higher than Devin’s 0.4%), and chore tasks showed 4.0% (4x higher than test tasks’ 1.0%).
4.2. RQ2: PR-MCI Taxonomy
Figure 3. Distribution of PR-MCI categories overall, by coding agent, and by task type.
Figure 3 shows the overall taxonomy, by agent, and by task type. Overall, the most common PR-MCI type was Phantom Changes (45.4%), where PR descriptions claim unimplemented changes. This was followed by Scope Understated (22.0%), where descriptions omit significant changes, and Placeholder/Incomplete (18.8%), which used generic or boilerplate text [^2] Examples of different PR-MCI types are provided in our replication package..
For agent-specific patterns, GitHub Copilot was dominated by Phantom Changes (74.0%), while Cursor (52.8%) and Devin (41.0%) showed Scope Understated as the primary issue. OpenAI Codex showed a balanced distribution with Placeholder/Incomplete (39.8%) and Scope Understated (41.9%). Task-specific patterns showed that Test (55.0%) and Bug Fix (53.5%) PRs were most prone to Phantom Changes, while Refactor PRs showed Scope Understated (40.5%) as the dominant issue.
RQ2: Phantom Changes was the dominant PR-MCI type (45.4%), followed by Scope Understated (22.0%) and Placeholder/Incomplete (18.8%), suggesting that AI agents frequently claim unimplemented changes or omit significant changes.
4.3. RQ3: Association of PR-MCI with Review and Outcome
Table 2. Review and Outcome Metrics by MCI Level. Highlighted values indicate statistically significant differences between High-MCI and Low-MCI PRs with non-negligible effect sizes. See §3 for statistical test details.
Metric Low-MCI High-MCI $p$-value Effect Size # (Low) # (High) Acceptance Rate (%) 80.0 28.3 <<0.001 Small 22,841 406 Mean Time to Merge (h) 16.0 55.8 <<0.001 Small 18,282 115 Mean # Review 0.74 0.42 0.039 Negligible 22,841 406 Mean # Comment 0.64 0.37 0.012 Negligible 22,841 406 Acceptance Rate by Agent (%) Claude Code 69.3 50.0 1.000 Negligible 218 2 GitHub Copilot 59.4 3.4 <<0.001 Medium 2,441 234 Cursor 73.3 67.7 0.638 Negligible 651 31 Devin 58.3 45.5 0.580 Negligible 2,880 11 OpenAI Codex 87.2 62.5 <<0.001 Negligible 16,651 128 Acceptance Rate by Task Type (%) Chore 80.2 45.8 <<0.001 Small 576 24 Documentation 89.8 34.5 <<0.001 Small 2,786 29 Feature 80.1 27.7 <<0.001 Small 10,782 166 Bug Fix 73.6 17.7 <<0.001 Small 5,206 113 Refactor 78.1 42.6 <<0.001 Small 1,499 54 Test 84.3 25.0 <<0.001 Small 1,992 20 Time to Merge (h) by Agent (merged PRs only) Claude Code 31.4 6.1 0.817 Negligible 149 1 GitHub Copilot 70.2 24.0 0.163 Small 1,420 8 Cursor 23.1 28.9 0.333 Negligible 477 21 Devin 21.8 14.4 0.976 Negligible 1,669 5 OpenAI Codex 5.7 38.6 <<0.001 Medium 14,509 77 Time to Merge (h) by Task Type (merged PRs only) Chore 16.5 81.4 0.181 Small 460 11 Documentation 8.9 1.1 0.140 Small 2,493 10 Feature 11.3 16.4 <<0.001 Medium 8,617 44 Bug Fix 20.6 34.4 <<0.001 Medium 3,812 20 Refactor 15.5 51.2 0.012 Small 1,165 22 Test 6.4 80.6 0.825 Negligible 1,677 5
Table 2 reports summary statistics for key metrics with statistical tests and breakdowns by agent and task type. Notably, high-MCI PRs were associated with a significantly lower acceptance rate: 28.3% vs. 80.0%, a difference of 51.7% ($p<0.001$). Effect size was small (Cramér’s V = 0.166), indicating a non-trivial practical difference. The relationship varied substantially by agent: for GitHub Copilot, the effect was largest (3.4% vs. 59.4%, 55.9 percentage point difference, $p<0.001$). The effect of high-MCI was even more obvious by task type, with statistically significant differences for all tasks (all $p<0.001$ with small effect sizes).
For merged PRs, high-MCI PRs took significantly longer to merge: 55.8 hours vs. 16.0 hours ($p<0.001$; Cliff’s $\delta=0.310$). The effect was most pronounced for OpenAI Codex (38.6 vs. 5.7 hours, $p<0.001$) and refactoring tasks (51.2 vs. 15.5 hours, $p=0.012$).
Further, regression analysis [^3] Detailed regression results are available in our replication package. showed that PR-MCI remained significantly associated with lower acceptance and longer merge times ($p<0.001$ for both) after controlling for code churn, files changed, task type, and agent, suggesting the association is not fully explained by code complexity or task characteristics alone.
[…]
5. Discussion
5.1. Actionable Insights for Developers
Our findings demonstrate that even when AI-generated code is acceptable, PR descriptions may misstate scope or claim phantom changes, confirming why human oversight remains necessary for Agentic-PRs (Watanabe et al., 2025). Therefore, we suggest:
For agent users: (1) verify that agents actually commit changes, (2) avoid template-based prompts, and (3) refine prompts to guide agents toward more complete descriptions.
For PR reviewers, we recommend three quick heuristic checks to identify PR-MCI without examining the diff in detail:
For PR reviewers: (1) verify the diff is not empty, (2) check if descriptions seem too brief for the amount of code changed, and (3) check for template markers (e.g., [WIP], [Failed], or generic phrases like ”I’m starting to work on it”).
5.2. Actionable Insights for AI Tool Builders
[…]
For AI tool builders: (1) implement automated PR-MCI verification mechanisms that can detect common types (e.g., Phantom Changes, Scope Understated, and Placeholder/Incomplete), and (2) integrate them into Agentic-PR generation pipelines to flag problematic descriptions before review, helping recover thousands of hours of wasted review effort.
5.3. Implications for SE 3.0 Research
The emerging “SE 3.0” vision positions AI as active teammates in the software development lifecycle (Li et al., 2025; Gröpler et al., 2025), where AI agents are expected not only to generate code, but also to communicate intent, justify changes, and support human decision-making in collaborative workflows. Our findings provide empirical evidence that current Agentic-PR descriptions may vary substantially in quality and can strongly affect review outcomes.
Building on this insight, future research could leverage our 974 manually annotated PRs with 432 partial/misaligned cases to:
For Researchers: (1) investigate why agent communication abilities vary, (2) develop standardized evaluation metrics for PR description quality, and (3) explore how Agentic-PR generation can be improved through better prompt engineering (Gong et al., 2025b; Huang et al., 2024), fine-tuning (Huang et al., 2025a; Gong et al., 2025c), or reinforcement learning (Gong, 2025; Türker et al., 2024; Thomas et al., 2025).
5.4. Threats to Validity
This study has several limitations that should be considered when interpreting the results. (1) Construct validity: Our PR-MCI metric is a heuristic proxy for semantic consistency and may miss subtle mismatches or penalize concise but accurate descriptions. (2) Threshold calibration: The decision threshold is calibrated on a validation set, which weakens absolute prevalence claims despite observed stability under cross-validation. (3) External validity: Dataset filtering choices prioritize scale over strict controls and may affect generalizability. (4) Sample size imbalance: Agent sample sizes are imbalanced (e.g., Claude Code has only 220 PRs), limiting the reliability of agent-specific conclusions for smaller samples. (5) Taxonomy completeness: The taxonomy is derived from a limited subset of PRs (432 partial/misaligned cases) and may not capture all inconsistency types. (6) Causal inference: Our analysis is observational and does not establish causal relationships between PR-MCI and review outcomes.
5.5. Ethical Considerations
This study analyzes only publicly available GitHub data from repositories under permissive licenses (MIT or Apache 2.0) and reported findings in aggregate to avoid identifying individuals or projects.
6. Conclusion
We used PR-MCI to quantify alignment between AI-generated PR descriptions and code changes, analyzing 23,247 PRs from five AI agents. Our findings reveal that unreliable PR descriptions (high PR-MCI) are associated with significantly lower PR acceptance rates and longer merge times, indicating that improving the consistency and accuracy of AI-generated PR descriptions is essential for trustworthy human-AI collaboration.
Data Availability. All data, scripts, results, extra analysis, and supplementary materials are available in our replication package.